# Maximising Histopathology Segmentation using Minimal Labels via Self-Supervision

### v1

@@ Abstract
Histopathology, the microscopic examination of tissue samples, is essential for disease diagnosis and prognosis. Accurate segmentation and identification of key regions in histopathology images are crucial for developing automated solutions. However, state-of-art deep learning segmentation methods like UNet require extensive labels, which is both costly and time-consuming, particularly when dealing with multiple stainings. To mitigate this, multi-stain segmentation methods such as MDS1 and UDAGAN have been developed, which reduce the need for labels by requiring only one (source) stain to be labelled. Nonetheless, obtaining source stain labels can still be challenging, and segmentation models fail when they are unavailable. This article shows that through self-supervised pre-training—including SimCLR, BYOL, and a novel approach, HR-CS-CO—the performance of these segmentation methods (UNet, MDS1, and UDAGAN) can be retained even with 95% fewer labels. Notably, with self-supervised pre-training and using only 5% labels, the performance drops are minimal: $5.9\%$ for UNet, $4.5\%$ for MDS1, and $6.2\%$ for UDAGAN, compared to their respective fully supervised counterparts (without pre-training, using 100% labels). The code is available from https://github.com/zeeshannisar/improve_kidney_glomeruli_segmentation [to be made public upon acceptance].
@@
journal: Pattern Recognition
@@ 1 Introduction
To overcome these challenges, various stain transfer based multi-stain segmentation methods have been developed that are trained using labels from only one (source) stain. The current state-of-the-art (SOTA) approaches in this field include: Multi-Domain Supervised 1 (MDS1), which allows the application of a segmentation model trained on a source stain to multiple target stains by translating them to source stain; and (b) Unsupervised Domain Augmentation using Generative Adversarial Networks (UDAGAN), an approach to train a single stain-invariant segmentation model on multiple stains but using labels from only the source stain. These methods achieve results comparable to the fully supervised UNet (as will be seen later in this study).
@@
While effective, these methods rely on substantial amounts of source stain labelled data, which can be hard to obtain. In parallel, the size of unlabelled medical imaging datasets are increasing, e.g.  the advent of whole slide imaging (WSI) scanners has facilitated the production of vast amounts of histopathological image data. The primary focus of this study is to show that this unlabelled data can counteract the effects of a lack of labelled data in segmentation tasks through self-supervised learning (SSL).
@@
SSL learns representations using unlabelled data by designing a pretext task [3], this representation can then be used in various downstream tasks where labelled data is limited. Accordingly, this study focuses on using the state-of-art SSL methods, such as SimCLR [4], BYOL [5], and our proposed extension to CS-CO [6], called HR-CS-CO, that overcomes its stain-specific limitation. Many SSL studies focus on classification, however there is a lack of studies focusing on segmentation, particularly multi-stain segmentation. Therefore, the downstream segmentation tasks studied herein are: (a) single-stain segmentation using UNet; (b) multi-stain segmentation using MDS1; and (c) stain-invariant segmentation, as multi-stain, using UDAGAN.
@@
Notably, this article will show that when pre-trained with SSL and fine-tuning with few labels, performance comparable to fully (100%) supervised models can be obtained with all the above-mentioned tasks. For instance, five UNets fine-tuned with only $\sim$30 labels, each from one of five different stains, achieve an average F1 score of $0.810$ across all stains, compared to $0.869$ for their fully supervised counterparts (each trained with $600$ labels). Similarly, fine-tuned MDS1 and UDAGAN models achieve average F1 scores of $0.744$ and $0.765$ (respectively), across five stains, using only $30$ labels from one (source) stain, whereas their respective baselines trained using $600$ labels achieve $0.789$ and $0.827$ respectively. These findings demonstrate that self-supervised fine-tuned models can approach the performance of fully supervised models, while significantly reducing label requirements.
@@
    Conduct the first comprehensive analysis of SSL’s impact on multi-stain and stain-invariant segmentation, using the use case of kidney glomeruli segmentation to deduce the most appropriate self-supervised approach.
@@
    Reduce the annotation requirement of MDS1 for multi-stain segmentation and UDAGAN for stain-invariant segmentation by more than 95%, meaning a stain-invariant model can be trained from as few as $\sim$30 labelled glomeruli.
@@
The rest of the article is organised as follows: Section 2 reviews the literature on SSL; Section 3 provides the architectural details of the afore-mentioned self-supervised pre-training methods and downstream segmentation tasks; Section 4 provides an overview of the dataset and explains the training details; Section 5 evaluates the effectiveness of self-supervised pre-training and provide a detailed comparison to their baseline results; Section 6 discusses these results alongside the limitations and potential applications of our study; and finally, Section 7 concludes the article.
@@ 3.1 Self-Supervised Pre-training
This study uses three approaches to self-supervised pre-training: SimCLR [4], BYOL [5] and CS-CO [6]. SimCLR is selected due to its widespread adoption in histopathology related downstream tasks [11, 12]. While SimCLR’s performance heavily depends on augmentation and the number of negative pairs, demanding huge computing resources [4, 11], recent methods have shown that negative-pairs are not essential for contrastive learning [5, 18, 19]. One such method is BYOL [5]. The original architectures, as proposed by the authors of SimCLR and BYOL, were used in this study, with details provided in Section 3.1.1 and 3.1.2 respectively. Finally, to assess the benefits of hybrid SSL strategies, a novel extension to CS-CO [6] is included. This extension (presented in Section 3.1.3) overcomes its stain specific formulation (it was proposed for the H&E stain) and is called HR-CS-CO.
3.1.1 SimCLR
Fig. 1: Overview of the SimCLR architecture inspired by [4, 12].
SimCLR [4] learns representations by maximising the agreement between two augmented views of the same image via contrastive loss in the latent space. As shown in Fig. 1, the framework starts with a probabilistic data augmentation module $f_{aug}$ that generates two positively correlated views, $x_{i}$ and $x_{j}$, of a given data sample $x$. A set of base augmentations [4] are adopted (using the Albumentations library [20] and parameters from [11]), including random cropping and resizing with a large scale range of ($0.2-1.0$), flipping, grey-scale, Gaussian blur, and random colour distortions. Based on the findings of Stacke et al. [11], two additional augmentations were used, grid distort and grid shuffle, which have demonstrated their effectiveness for histopathology related applications. The augmented views, $x_{i}$ and $x_{j}$ are then transformed into their corresponding representations, $h_{i}$ and $h_{j}$ by employing a convolutional neural network (CNN) base encoder $f_{\theta}$, where $\theta$ is the weight parameters.
where $\tau$ is the temperature parameter that weights different samples and facilitates learning from hard negative samples and $\mathds{1}$ is the indicator function. The term $sim(z_{i},z_{j})=z_{i}^{\top}z_{j}/\|z_{i}\|\|z_{j}\|$ represents the dot product between $\ell_{2}$ normalised $z_{i}$ and $z_{j}$, which corresponds to the cosine similarity. This loss functions aims to maximise the agreement between positive pairs of augmented images, while minimising it for other images in the same batch (negative pairs). In each training step with a batchsize of $2N$, each augmented image has one positive and $2(N-1)$ negative pairs.
3.1.2 BYOL
Fig. 2: Overview of BYOL architecture inspired by Grill et al. [5].
$\xi\leftarrow\tau\xi+(1-\tau)\theta.$ (2)
The loss $\mathcal{L}_{\theta,\xi}$ is made symmetrical by separately feeding $x_{j}$ to the Online network and $x_{i}$ to the Target network. This allows the computation of another loss function ~$\tilde{\mathcal{L}}_{\theta,\xi}$. During each training step, a stochastic optimisation step is performed to minimise $\mathcal{L}^{BYOL}_{\theta,\xi}=\mathcal{L}_{\theta,\xi}+\tilde{\mathcal{L}}_{\theta,\xi}$ with respect to $\theta$ only, while $\xi$ remains unaffected by applying a stop-gradient ($sg$) as shown in Fig. 2.
@@ 3.1.3 CS-CO
Nevertheless, CS-CO has certain limitations, restricting its broader applicability. For example, histopathological images often use different staining protocols and reagents to highlight different tissue structures (e.g. PAS, Jones H&E, Sirius Red, CD68, and CD34, as used in this study). The stain separation method integral to CS-CO struggles with ImmunoHistochemical (IHC) stains [21]. Particularly, it fails to accurately extract the individual $H_{ch}$ and $DAB_{ch}$ (Diaminobenzidine) from CD68. Furthermore, in some cases, histopathological stains contain more than two dyes, e.g. Jones H&E, where CS-CO’s stain-separation approach would yield three separate channels—$J_{ch}$ (Jones), $H_{ch}$, and $E_{ch}$—which cannot be handled in CS-CO’s architecture.
@@
where $\bar{w}^{\prime}$, $\bar{w}^{\prime\prime}$, $\bar{z}^{\prime}$, and $\bar{z}^{\prime}$ are the $\ell_{2}$ normalised versions of $w^{\prime}$, $w^{\prime\prime}$, $z^{\prime}$, and $z^{\prime\prime}$, respectively. This encourages $w^{\prime}$ and $w^{\prime\prime}$ to be similar to $z^{\prime\prime}$ and $z^{\prime}$ (respectively). Before computing the loss, the stop-gradient ($sg$) is applied to $z^{\prime}$ and $z^{\prime\prime}$. This $sg$ step introduces a necessary asymmetry in gradient flow, which, when combined with the symmetrised loss, allows robust and diverse feature learning. During contrastive learning, the frozen pre-trained HR decoder ($\psi$) continues to use the outputs of the HR encoder ($\phi$) for image reconstruction. To avoid collapse, the HR encoder must maintain the necessary information for image reconstruction, by satisfying Equation 4. As a result, the total loss is formulated as $\mathcal{L}_{csco}=\mathcal{L}_{cs}+\gamma\mathcal{L}_{co}$, where $\gamma$ is the weight coefficient.
@@ 3.2 Downstream Segmentation Tasks
Once a model has been pre-trained using one of the self-supervised pre-training methods mentioned above, it can be fine-tuned for various downstream tasks. The primary aim of this study is to investigate the use of these pre-training methods to reduce the need of labels for several segmentation tasks, (a) UNet [2], (b) MDS1 [23], and (c) UDAGAN [24], as shown in Fig. 6. The original architectures as proposed by the authors were used, with details provided below. Additional details on combining these with pre-training are presented in Section 4.2.
@@ 3.2.1 UNet
3.2.2 MDS1 and UDAGAN
Fig. 8: Overview of MDS1 and UDAGAN architecture.
These are stain transfer based multi-stain segmentation approaches, that use labels from the source ($S$) stain, which is PAS in our case, and eliminate the need for labels in the target ($T$) stains (herein these are Jones H&E, Sirius Red, CD68, CD34). Stain transfer is achieved using a cycle-consistent generative adversarial network (CycleGAN) [26]. For both MDS1 and UDAGAN, a separate CycleGAN model ($M^{S\leftrightarrow T}_{\text{cycGAN}}$) is trained to translate between the $S$ and each $T$ stain, i.e. from $S\rightarrow T$ and $T\rightarrow S$, see Fig. 8. The architectural descriptions for CycleGAN are detailed in [26].
In MDS1 [23], see Fig. 8, a UNet segmentation model ($M^{S}_{\text{unet}}$), as detailed above, is trained using the source stain and its labels. Then, during testing, the target stain is translated to match the source, referred to as the ‘fake S’, and $M^{S}_{\text{unet}}$ is applied on ‘fake S’ to obtain segmentation masks.
In contrast, UDAGAN [24] combines stain augmentation and adaptation to create a stain-invariant model that can be directly applied to multiple stains without test-time translation. Specifically, CycleGAN models ($M^{S\rightarrow T}_{\text{cycGAN}}$) are used to augment the labelled source stain by randomly translating images to one of the target stains. Since translation does not change the image’s structure, see Fig. 8, it’s ground truth is still valid. As such, several annotated samples from all available stains are presented to the UNet, resulting in a single stain-invariant model able to segment several target stains.
@@ 4.1 Dataset
The findings presented herein were evaluated for kidney glomeruli segmentation using a private histopathology dataset provided by the Hannover Medical School. Tissue samples were collected from $10$ patients who had tumor nephrectomy due to renal carcinoma. The renal tissue was selected as distant as possible from the tumors to represent largely normal renal glomeruli. However, certain samples exhibited varying degrees of pathological modifications, such as complete or partial displacement of functional tissue by fibrotic changes (“scerosis”) indicating normal age-related changes or the renal effects of general cardiovascular comorbidity (e.g. cardial arrhythmia, hypertension, arteriosclerosis). Using the Ventana Benchmark Ultra automated staining tool, the paraffin-embedded samples were sliced into 3µm thick sections and stained with either Jones H&E, PAS, Sirius Red, and two immunohistochemistry markers (CD34, and CD68). An Aperio AT2 scanner was used to capture whole slide pictures at $40$ magnification (a resolution of $0.253$ m / pixel), resulting in WSIs ranging from 37k×$\times$25k to 107k×$\times$86k pixels. Pathologists annotated and verified all of the glomeruli in each whole slide image (WSI) by labelling them with Cytomine [27]. The dataset was split into $4$ training, $2$ validation, and $4$ test patients. The level-of-detail used is $1$ (corresponding to 20$\times$ magnification) with a patch size of $508\times 508$ pixels, as to contain a glomeruli with its surrounding area.
@@ 4.1.2 Downstream Segmentation Tasks
MDS1 & UDAGAN –
To train the respective CycleGAN models, $5000$ patches are randomly extracted from the training WSIs of each stain (PAS being the source, and the rest the targets). Additionally, when fine-tuning the pre-trained MDS1 and UDAGAN models, the same splits as used for the UNet were used, but only those of the PAS stain, see Table 4.1 ($1^{\text{st}}$ column).
@@ 4.2.2 Downstream Segmentation Tasks
As is common, the validation loss (see Section 4.1.2) is used to select the final model [11, 6, 12]. This is, however, a hindrance in histopathology (and medical imaging in general) since it requires additional labels. Thus, additional experiments were performed without a validation set by selecting the best model based on the training loss. The training details for downstream tasks (UNet, MDS1, and UDAGAN) and the strategy to combine the pre-trained features from each of the self-supervised pre-trained method are detailed below.
@@
MDS1:
MDS1 requires both a UNet and CycleGAN network to be trained. The UNet is trained as described above. The key difference being that it uses only labelled splits of the source stain (PAS). For CycleGAN, the network architecture and loss weights ($w_{cyc}=10,w_{id}=5$) are taken from the original paper [26]. To deal with large patch sizes (i.e. above $256\times 256$ pixels), a translation network with nine ResNet blocks is employed. The model is trained for $50$ epochs, with a LR of $0.0002$ using the Adam optimiser, and a batch size of $1$. Starting from the $25^{\text{th}}$ epoch, the LR is linearly decayed to 00, and the weights for cycle-consistency ($w_{cyc}$) and identity ($w_{id}$) are halved. In all experiments, the translation model from the final epoch is used. To reduce model oscillation, Shrivastava et al. [29]’s strategy is adopted that updates the discriminator using the $50$ previously generated samples. To account for random variations, the CycleGAN network was trained three times per target stain, and five UNets were trained per CycleGAN (resulting in 15 repetitions for each split of the source stain labels).
@@
The training of UDAGAN is similar to MDS1, incorporating both UNet and CycleGAN networks. Specifically, in the first step, CycleGAN network is trained for each target stain, enabling the translation from $S\rightarrow T$. Subsequently, a training patch from $S$ stain is translated into randomly selected $T$ stain (with a probability of $\frac{N-1}{N}$, where $N$ is the number of stains) using the pre-trained CycleGAN network(s). Thus, all available stains (including the $S$ stain) are presented to the UNet model with equal probability, $\frac{1}{N}$, forcing the network to learn stain invariant features. The training details for CycleGAN and UNet are the same as those outlined in MDS1. One UDAGAN model is trained for each split of the $S$ stain labels and repeated five times.

### v2

# Maximising Kidney Glomeruli Segmentation using Minimal Labels via Self-Supervision

@@ Abstract
Histopathology, the microscopic examination of tissue samples, is essential for disease diagnosis and prognosis. Accurate segmentation and identification of key regions in histopathology images are crucial for developing automated solutions. However, state-of-art deep learning segmentation methods like UNet require extensive labels, which is both costly and time-consuming, particularly when dealing with multiple stainings. To mitigate this, various multi-stain segmentation methods such as UDAGAN have been developed, which reduce the need for labels by requiring only one (source) stain to be labelled. Nonetheless, obtaining source stain labels can still be challenging, and segmentation models fail when they are unavailable. This article shows that through self-supervised pre-training—including SimCLR, BYOL, and a novel approach, HR-CS-CO—the performance of these segmentation methods (UNet, and UDAGAN) can be retained even with 95% fewer labels. Notably, with self-supervised pre-training and using only 5% labels, the performance drops are minimal: $5.9\%$ for UNet and $6.2\%$ for UDAGAN, compared to their respective fully supervised counterparts (without pre-training, using 100% labels). Furthermore, these findings are shown to generalise beyond their training distribution to public benchmark datasets. Implementations and models are available at https://github.com/zeeshannisar/maximising-kidney-glomeruli-segmentation-using-limited-labels.
@@
journal: Computers in Biology and Medicine
@@ 1 Introduction
To overcome these challenges, various stain transfer based multi-stain segmentation methods [3, 4, 5, 6] have been developed. One of the most prominent approach is Unsupervised Domain Augmentation using Generative Adversarial Networks (UDAGAN) [4]. This method enables the training of a single stain-invariant segmentation model on multiple stains using labels from only the source stain. These methods show potential to achieve results comparable to the fully supervised UNet (as will be seen later in this study).
@@
While effective at reducing the amount of labelled data to only a single stain, such methods still rely on substantial amounts of source stain labelled data, which can be hard to obtain. In parallel, the size of unlabelled medical imaging datasets are increasing, e.g.  the advent of whole slide imaging (WSI) scanners has facilitated the production of vast amounts of histopathological image data. The primary focus of this study is to show that this unlabelled data can counteract the effects of a lack of labelled data in segmentation tasks through self-supervised learning (SSL).
@@
By learning representations using unlabelled data using a pretext task [7], a SSL representation can then be used in various downstream tasks where labelled data is limited. Accordingly, this study focuses on using the state-of-art SSL methods, such as SimCLR [8], BYOL [9], and our proposed extension to CS-CO [10], called HR-CS-CO, that overcomes its stain-specific limitation. Many SSL studies focus on classification, however there is a lack of studies focusing on segmentation, particularly multi-stain segmentation. Therefore, the downstream segmentation tasks studied herein are: (a) single-stain segmentation using UNet; and (b) stain invariant based multi-stain segmentation using UDAGAN.
@@
Notably, this article will show that when pre-trained with SSL and fine-tuning with few labels, performance comparable to fully (100%) supervised models can be obtained in both of the above-mentioned tasks. For instance, five UNets fine-tuned with only 5% labels, each from one of five different stains, achieve an average F1 score of 0.810 across all stains, compared to 0.869 for their fully supervised counterparts (each trained with 100% labels). Similarly, fine-tuned UDAGAN models achieve average F1 scores of 0.765, across five stains, using only 5% labels from one (source) stain, whereas the respective baselines trained using 100% labels achieve 0.827. These findings demonstrate that self-supervised fine-tuned models can approach the performance of fully supervised models, while significantly reducing label requirements by up to 95%.
Additionally, our generalisation study reveals that the benefits of pre-trained models extend to diverse datasets and domains. For instance, on the HuBMAP Kaggle competition dataset, fine-tuned UNet models with just 5% labels achieve F1 score of 0.816, compared to 0.947 for fully supervised baseline and 0.951 for the competition winning model (trained with 100% labels). Similarly, on the KPIs Challenge dataset, fine-tuned UNets reach an F1 score of 0.795 using only 5% labels, outperforming the corresponding fully supervised baseline model (0.786) and approaching the top-performing models score of 0.886. These findings further highlight the effectiveness of these pre-trained models in learning robust and transferable features that enable competitive performance on external datasets even under extreme annotation scarcity.
@@
    Conduct the first comprehensive analysis of SSL’s impact on multi-stain histopathology segmentation, using the use case of kidney glomeruli segmentation to deduce the most appropriate self-supervised approach.
@@
    Reduce the annotation requirement of UDAGAN for stain-invariant based multi-stain segmentation by more than 95%.
@@
The rest of the article is organised as follows: Section 2 reviews the literature on SSL; Section 3 provides the architectural details of the aforementioned self-supervised pre-training methods and downstream segmentation tasks; Section 4 provides an overview of the dataset and explains the training details; Section 5 evaluates the effectiveness of self-supervised pre-training and provide a detailed comparison to their baseline results; Section 6 examines how well this pre-training generalises across several external datasets; Section 7 discusses these results alongside the limitations and potential applications of our study; and finally, Section 8 concludes the article.
@@ 3.1 Self-supervised Pre-training
This study uses three approaches to self-supervised pre-training: SimCLR [8], BYOL [9] and CS-CO [10]. SimCLR is selected due to its widespread adoption in histopathology related downstream tasks [15, 16]. While SimCLR’s performance heavily depends on augmentation and the number of negative pairs, demanding huge computing resources [8, 15], recent methods have shown that negative-pairs are not essential for contrastive learning [9, 22, 23]. One such method is BYOL [9]. The original architectures, as proposed by the authors of SimCLR and BYOL, were used in this study, with details provided in A.1.1 and A.1.2 respectively. Finally, to assess the benefits of hybrid SSL strategies, a novel extension to CS-CO [10] is included. This extension (presented in Section 3.1.1) overcomes its stain specific formulation (it was proposed for the H&E stain) and is called HR-CS-CO.
@@ 3.1.1 CS-CO
[…]
@@
where $\bar{w}^{\prime}$, $\bar{w}^{\prime\prime}$, $\bar{z}^{\prime}$, and $\bar{z}^{\prime}$ are the $\ell_{2}$ normalised versions of $^{\prime}, w′′w^{\prime\prime}$, $z^{\prime}$, and $z^{\prime\prime}$, respectively. This encourages $w^{\prime}$ and $w^{\prime\prime}$ to be similar to $z^{\prime\prime}$ and $z^{\prime}$ (respectively). Before computing the loss, the stop-gradient ($sg$) is applied to $z^{\prime}$ and $z^{\prime\prime}$. This $sg$ step introduces a necessary asymmetry in gradient flow, which, when combined with the symmetrised loss, allows robust and diverse feature learning. During contrastive learning, the frozen pre-trained HR decoder ($\psi$) continues to use the outputs of the HR encoder ($\phi$) for image reconstruction. To avoid collapse, the HR encoder must maintain the necessary information for image reconstruction, by satisfying Equation 1. As a result, the total loss is formulated as $\mathcal{L}_{csco}=\mathcal{L}_{cs}+\gamma\mathcal{L}_{co}$, where $\gamma$ is the weight coefficient.
@@ 3.2 Downstream Segmentation Tasks
Once a model has been pre-trained using one of the self-supervised pre-training methods mentioned above, it can be fine-tuned for various downstream tasks. The primary aim of this study is to investigate the use of these pre-training methods to reduce the need of labels for several segmentation tasks, such as UNet [2], and UDAGAN [4], as shown in Fig. 4. The original architectures as proposed by the authors were used, with details provided below. Additional details on combining these with pre-training are presented in Section 4.2.
@@ 3.2.1 UNet
3.2.2 UDAGAN
UDAGAN [4] is an advanced stain transfer based multi-stain segmentation approach, that use labels from the source ($S$) stain, which is PAS in our case, and eliminate the need for labels in the target ($T$) stains (herein these are Jones H&E, Sirius Red, CD68, CD34). Stain transfer is achieved using a cycle-consistent generative adversarial network (CycleGAN) [27]. Specifically, a separate CycleGAN model ($M^{S\leftrightarrow T}_{\text{cycGAN}}$) is trained to translate between the $S$ and each $T$ stain, i.e. from $S\rightarrow T$ and $T\rightarrow S$, see Fig. 6. The architectural descriptions for CycleGAN are detailed in [27].
Fig. 6: Overview of UDAGAN architecture.
Once the CycleGAN models ($M^{S\rightarrow T}_{\text{cycGAN}}$) are trained, they are used to augment the labelled source stain by randomly translating images to one of the target stains. Since translation does not change the image’s structure, see Fig. 6, it’s ground truth is still valid. As such, samples from all available stains, but that are annotated in the source stain, are presented to the UNet, resulting in a single stain-invariant segmentation model—UDAGAN—that can be directly applied to multiple stains.
@@ 4.1 Dataset
[…]
@@ 4.1.2 Downstream Segmentation Tasks
UDAGAN –
To train the respective CycleGAN models, $5000$ patches are randomly extracted from the training WSIs of each stain (PAS being the source, and the rest the targets). Additionally, when fine-tuning the pre-trained UDAGAN models, the same splits as used for the UNet are preserved, but only the samples from the PAS stain are used, see Table 4.1 ($1^{\text{st}}$ column).
@@ 4.2.2 Downstream Segmentation Tasks
As is common, the validation loss (see Section 4.1.2) is used to select the final model [15, 10, 16]. This is, however, a hindrance in histopathology (and medical imaging in general) since it requires additional labels. Thus, additional experiments were performed without a validation set by selecting the best model based on the training loss. The training details for both downstream tasks (UNet, and UDAGAN) and the strategy to combine the pre-trained features from each of the self-supervised pre-trained method are detailed below.
@@
Training the UDAGAN involves both UNet and CycleGAN networks. Specifically, in the first step, a CycleGAN network is trained for each target stain, enabling the translation from $S\rightarrow T$. To train these CycleGAN models, the network architecture and loss weights ($w_{cyc}=10,w_{id}=5$) are taken from the original paper [27]. To deal with large patch sizes (i.e. above $256\times 256$ pixels), a translation network with nine ResNet blocks is employed. The model is trained for $50$ epochs, with a LR of $0.0002$ using the Adam optimiser, and a batch size of $1$. Starting from the $25^{\text{th}}$ epoch, the LR is linearly decayed to 0, and the weights for cycle-consistency ($w_{cyc}$) and identity ($w_{id}$) are halved. In all experiments, the translation model from the final epoch is used. To reduce model oscillation, Shrivastava et al. [30]’s strategy is adopted that updates the discriminator using the $50$ previously generated samples. After training CycleGAN, a training patch from $S$ stain is translated to a randomly selected $T$ stain (with a probability of $\frac{N-1}{N}$, where $N$ is the number of stains) using the pre-trained CycleGAN network(s). Thus, all available stains (including the $S$ stain) are presented to the UNet model with equal probability, $\frac{1}{N}$, forcing the network to learn stain invariant features. One UDAGAN model is trained for each split of the $S$ stain labels and repeated five times.